# Claim Verification in the Age of Large Language Models: A Survey

### v1

@@ Introduction
Retrieved evidence may be irrelevant, which is a challenge for LLMs, as they may not be trained to ignore such evidence. The lack of robustness to the noise can cause the LLM to produce misinformation and incorrect verification. Recent research on open-domain question answering shows that external knowledge relevant to the task can aid the model performance, however, irrelevant context can also lead the model to make inaccurate predictions (Petroni et al. 2020; Shi et al. 2023; Li et al. 2023a; Yu et al. 2023). For the fact-verification task, recent works have proposed techniques to identify the most relevant context while aiding the veracity label prediction and thus improving the overall system performance (Wang et al. 2023c; Yoran et al. 2024; Xia et al. 2024). However, more work on the topic is required to verify the effectiveness of this design approach across multiple fact verification domains.
@@
Fact-verification approaches reliance on retrieved evidence can cause knowledge conflicts in LLM-based approaches where retrieved evidence as the internal parameters of the pre-trained LLM may conflict with the external knowledge. This causes the LLM to ignore the retrieved evidence and produces hallucinations. Xu et al. (2024a) provides an in-depth analysis of this scenario. Works by (Li et al. 2023a; Neeman et al. 2023; Mallen et al. 2023; Longpre et al. 2021; Chen, Zhang, and Choi 2022) introduce approaches to avoid knowledge conflicts and mitigate hallucinations for question-answering. Expanding this work to fact-verification, especially to the approaches that use LLMs, is vital for an effective verification process.
@@
Most automated claim verification approaches rely on English datasets. Furthermore, there are limited multilingual fact-verification datasets (Gupta and Srikumar 2021; Kazemi et al. 2022; Pikuliak et al. 2023; Singh et al. 2023). This hinders the development of approaches for multilingual fact-verification, which achieve the best performance when trained on language-specific datasets (Panchendrarajan and Zubiaga 2024).
Conclusion
We presented a survey on LLM approaches to claim verification. To the best of our knowledge, this is the first claim verification survey to focus exclusively on LLM approaches thus filling an important gap in the literature. We have described each of the sub-tasks of the typical claim verification pipeline and discussed various LLM-based approaches used in this task. Finally, we have also described publicly available English datasets providing important information to new and seasoned researchers on this topic.
Advances in LLM development will likely continue improving the quality of claim verification systems. We hope this survey motivates future research on this topic taking advantage of recently-proposed LLMs, RAG methods, etc. Claim verification is a vibrant research topic and we see multiple open research directions as described in the next sub-section.

### v2

@@ 6 Open Challenges
Retrieved evidence may be irrelevant, which is a challenge for LLMs, as they may not be trained to ignore such evidence. The lack of robustness to noise can lead an LLM to produce misinformation and incorrect verification. Recent research on open-domain question answering shows that external knowledge relevant to the task can improve model performance, however, irrelevant context can also lead to inaccurate predictions Petroni et al. (2020); Shi et al. (2023); Li et al. (2023a); Yu et al. (2023). For the fact-verification, recent work has proposed techniques to identify the most relevant context for improving veracity label prediction, thus improving the overall system performance Wang et al. (2023); Yoran et al. (2024); Xia et al. (2024). However, more work is required to verify the effectiveness across domains.
@@
The reliance of fact-verification approaches on retrieved evidence can cause knowledge conflicts in LLMs, where retrieved external evidence may conflict with the internal parameters of the pre-trained LLM. This causes the LLM to ignore the retrieved evidence and produces hallucinations Xu et al. (2024a). Approaches for avoiding knowledge conflicts have been introduced for question answering  Li et al. (2023a); Neeman et al. (2023); Mallen et al. (2023); Longpre et al. (2021); Chen et al. (2022a); Marjanović et al. (2024). Expanding this work to fact-verification, especially to the approaches that use LLMs, is vital for an effective verification process.
@@
Most automated claim verification approaches rely on English datasets. Furthermore, there are limited multilingual fact-verification datasets Gupta and Srikumar (2021); Kazemi et al. (2022); Pikuliak et al. (2023); Singh et al. (2023). This hinders the development of approaches for multilingual fact-verification, which achieve the best performance when trained on language-specific datasets Panchendrarajan and Zubiaga (2024).
7 Conclusion
We presented a survey on LLM approaches to claim verification. To the best of our knowledge, this is the first claim verification survey to focus exclusively on LLM approaches, thus filling an important gap in the literature. We have described the traditional claim verification pipeline including its component tasks and discussed various LLM-based approaches used in this task. Finally, we have also described publicly available English datasets providing important information to new and seasoned researchers on this topic.
Advances in LLM development will likely continue to improve the quality of claim verification systems. We hope this survey motivates future research on this topic taking advantage of recently-proposed LLMs, RAG methods, and others. Claim verification is a vibrant research topic and we envisage multiple open research directions such as handling irrelevant context, knowledge conflicts and multilingualism.